# Data-Driven High-Dimensional Statistical Inference with Generative Models

### v1

@@ 4 Example Application : Di-Higgs Measurement
We note that this number of background events is several orders of magnitude smaller than what can be expected in the HL-LHC dataset. Such a larger background sample would increase the statistics available to train the generative model in the sidebands, likely improving its quality and decreasing variability. However, this also means the signal injections we perform for testing our method, which consider signal strengths of S/B $\sim 10^{-2}-10^{-3}$ are all unrealistic for a true di-Higgs $bb\gamma\gamma$ analysis. It is likely that our background models would not be accurate at very small probabilities, of say $10^{-6}$, making it difficult to extract a signal with an S/B of this magnitude. However, one could imagine applying some additional selection criteria, perhaps using a multivariate classifier, to the data prior employing to HI-SIGMA, so as to enhance the expected signal fraction to the level of sensitivity of our generative modeling capabilities. Improvements to employed generative models could also increase the sensitivity of this method to smaller signal strengths HI-SIGMA. We leave these explorations for future work.
@@
SR :$\begin{split}\text{SR }&:m_{\gamma\gamma}\in[115,135)\text{ GeV}\\ \text{SB }&:m_{\gamma\gamma}\in[90,115)\cup[135,180)\text{ GeV}\\ \end{split}$ (10)
@@
Figure 1: Distributions of the signal (red) and background in the sideband region (blue) and signal regions (green). In the top plot the dashed red lines denote the boundaries of the signal region defined by $m_{\gamma\gamma}$. The distributions of background events in the signal and sideband regions are seen to be similar but with visible differences, which will be learned by the interpolation aspect of the HI-SIGMA approach.
@@ 4.1 Data Smearing
The amount of smearing applied to the data is therefore a very important choice, directly impacting the sensitivity of the final analysis. In this work, we found smearing each feature by 0.2 standard deviations allowed modeling of the distributions with high accuracy while resulting in only minor sensitivity loss. The resonance feature $m_{\gamma\gamma}$, which is not modeled by the normalizing flow, is not smeared. It is possible that with an improved machine learning model, less smearing, or perhaps no smearing, would be necessary. Alternatively perhaps a less simplistic smearing procedure, perhaps smearing some features and not others, or only for data near sharp boundaries could be employed to better mitigate sensitivity loss. As such considerations are very application dependent, we leave such explorations for future work.
A visualization of one feature distribution before and after smearing is shown in Fig. 2.
@@
4.2 Pre-processing and Model Architecture
@@
The generative models used to learn the signal and background conditional feature distributions, $P_{k}(\vec{x}^{\prime}|m)$, are normalizing flow using 5 rational quadratic spline layers MPRQAF2021 , each with 24 bins. BatchNorm batchnorm is used in between every flow layer. The flows are trained up to 100 epochs. Early stopping is triggered if no improvement is seen on the validation loss after 10 epochs. To account for training variability, five models are trained on the same dataset with different random seeds. The final model is taken as the ensemble of these 5 different models, averaging their log-probabilities to produce the final density estimate during evaluation.
@@ 4.4 Example Shape Uncertainty
To deploy of HI-SIGMA in any real measurement, it needs to be able to incorporate systematic uncertainties. To demonstrate that this is possible, we implement a toy shape systematic within this di-Higgs example following the method detailed in Sec. 3.2. We consider a $10\%$ relative uncertainty associated with the $p_{T}$ of the $H_{bb}$ candidate, such that for a given $\nu$ we re-scale $p^{bb}_{T}\to(1+0.1\nu)p^{bb}_{T}$ mimicking the effects of a jet energy scale uncertainty. This is done for both signal and background distributions, with both shifts parameterized by a single common shared nuisance parameter. We show the effects of the ±$\pm 1\sigma$ variations in Fig. 3.
Figure 3: Distributions of the $p_{T}^{bb}$ feature for $-\sigma$, nominal and $\sigma$ variations of the shape systematic. The other features are assumed to be unaffected.
For this simple uncertainty, the parameterization of the weights considered in 8 is too complex. We consider a simpler parameterization where $w_{k}(\vec{x}^{\prime},m,\nu)=1+\sum_{i=1}^{2}c_{ki}(\vec{x}^{\prime},m)\nu^{i}$ with no piecewise splitting. We performed a closure test to verify that the interpolation works as expected. Mock datasets were generated from the $\nu=0,\pm 1$ models, and successfully fit with the full model. We verified that the fitted signal strength and nuisance parameter agreed well with the true values, $(\hat{s},\hat{\nu})=(s_{\mathrm{true}},\nu_{\mathrm{true}})$.
@@ 4.5 Comparison Analysis Strategies
The next two analysis strategies employ DNN classifiers to discriminate signal versus background, similar to what is done in many current analyses. The first strategy employs a ’cut and fit’ approach. A DNN classifier is trained to distinguish signal versus background using the feature information, $\vec{x}^{\prime}$, but not the resonance feature $m_{\gamma\gamma}$. A cut is placed on the classifier score to reduce the background and the $m_{\gamma\gamma}$ distribution of the remaining events is fit to extract the signal. The cut value was chosen to maximize the signal significance, approximated as $S/\sqrt{B}$. Using a cut-based approach such as this is sub-optimal, as some signal events are excluded from the final analysis, and one throws away any information on signal versus background discrimination of the remaining events. Often, multiple exclusive categories based on different classifier score thresholds defined simultaneously fit, so as to mitigate this sensitivity loss. This strategy is employed in recent ATLAS and CMS analyses of di-Higgs $bb\gamma\gamma$ CMS_bbgg ; ATLAS_bbgg ; ATLAS_bbgg_run3 . However we stick with a single category here for simplicity.
@@
The third alternate analysis method uses a classifier trained using the full event information, the features and the resonance variable, to define an optimal summary statistic. A 1D histogram of the distribution of classifier scores in data is then fit to the sum of signal and background templates to extract the signal strength. Although the classifier will learn a monotonic rescaling of the likelihood ratio, the optimal observable for signal versus background discrimination, binning will generally result in some loss of information. In our example we choose quite a large number of bins, 50, to mitigate this sensitivity loss.
@@
Such an analysis requires an accurate model of the background classifier score shape. Because the classifier score distribution has complex multivariate correlations with all the input features, estimating the background shape using data-driven methods can be difficult and methods are usually analysis-specific. For purposes of sensitivity comparison, in this study we consider an idealized setting where the signal and background classifier score shapes are known exactly. This allows a comparison with the HI-SIGMA sensitivity in an idealized setting in which systematic uncertainties are not assessed on either method. However, it should be noted that the HI-SIGMA method automatically includes a data-driven background estimate, whereas for the classifier score analysis such an estimate may be quite difficult. Moreover, we should note that the optimality of the DNN-based analyses depends on the accuracy of the background simulations used for training, which is the very issue that motivates HI-SIGMA. In this sense, since we are assuming no background mismodelling when comparing methodologies, we aim to showcase how HI-SIGMA yields slightly improved performance without relying on possibly inaccurate simulations.
@@
4.6 Results
The different analysis methods are evaluated on 5 independent data samples, each containing 40,000 events. For each data sample 200 signal events are injected, leading to an approximate significance with the mass-only fit of $\sim 3\sigma$. The different analysis methods are then applied to these datasets and compared. Evaluations with different signal injection sizes were performed and found to give similar results. Examples of the mass only fit, classifier cut and fit, and binned DNN score fit are shown in Fig. 4.
Figure 4: Examples of the mass only fit (top), cut+fit approach (lower left) and classifier score fit (lower right).
The HI-SIGMA fit across the multi-dimensional space can also be visualized. To do this, we sample events from the signal and background models to construct ‘templates’ of our signal and background distributions. These templates are then scaled by the best fit normalizations and compared to the observed data events. This comparison can be done in any way that is desired, and the simplest choice is to consider 1D projections of each feature, which we visualize as histograms. Examples are shown in Fig. 5. We stress that the binning is for visualization purposes only, as the actual model and fit are unbinned. For each feature one can see an excess of data events above the background model which is well-fit by the signal template.
@@ 4.6 Results
While the visualization in Fig. 5 is useful, these 1D projections do not ensure the data is being well described across all regions of the multidimensional space. Fortunately, because HI-SIGMA includes a full multi-dimensional model of the signal and background densities, one can check the fit agreement in sub-regions of the space easily. We consider a check of the most signal-sensitive region of the space, defined by placing a cut on the DNN classifier score $>0.9$ and restricting $m_{\gamma\gamma}$ to the signal region, 115–135 GeV. Placing these cuts on the events generated from our model as well as the data allows us to check modeling in this sub-region of the space. Such a comparison is shown in Fig. 6. We see that in this sub-region the data is well described by the model and signal contribution is much more visible. We stress that such a visualization can be performed for any sub-region of the feature space $\vec{x}^{\prime}$, but cannot be done for any region defined by an observable outside the feature space of the model (e.g. checking modeling in different regions of $\eta$ for our example). Although this method is helpful in assessing the quality of the fit in different regions of the multidimensional space, a full assessment of the goodness-of-fit of an unbinned high-dimensional fit would require further study Williams:2010vh .
Figure 6: Visualization of a subspace of the HI-SIGMA fit, the most signal-sensitive region, defined by a requirement on the supervised classifier score $>0.9$.
The average uncertainty on the signal strength, defined as $\frac{s}{s_{\mathrm{true}}}$, from the fit to the 5 independent datasets is summarized in Fig. 7. The HI-SIGMA approach is shown to give the best sensitivity, providing a $\sim 40\%$ improvement over the cut+fit approach and a $\sim 10\%$ improvement over the binned classifier score fit. We stress that this is an ‘idealized’ comparison, in which no systematic uncertainties have been included for any of the analysis strategies, and strategies for realistic estimates of the background in the classifier score fit have not been considered. The different analysis strategies would necessitate different systematic uncertainties, which may change their relative sensitivities. However, the improvements seen in this study nevertheless demonstrate the potential of the unbinned HI-SIGMA approach.

### v2

@@ 4 Example Application : Di-Higgs Measurement
We note that this number of background events is several orders of magnitude smaller than what can be expected in the HL-LHC dataset. Such a larger background sample would increase the statistics available to train the generative model in the sidebands, likely improving its quality and decreasing variability. However, this also means the signal injections we perform for testing our method, which consider signal strengths of S/B $\sim 10^{-2}-10^{-3}$ are all unrealistic for a true di-Higgs $bb\gamma\gamma$ analysis. It is likely that our background models would not be accurate at very small probabilities, of say $10^{-6}$, making it difficult to extract a signal with an S/B of this magnitude. However, one could imagine applying some additional selection criteria, perhaps using a multivariate classifier, to the data prior employing to HI-SIGMA, so as to enhance the expected signal fraction to the level of sensitivity of our generative modeling capabilities. Improvements to employed generative models could also increase the sensitivity of this method to smaller signal strengths. We leave these explorations for future work.
@@
[…]
@@
Figure 1: Distributions of the signal (red) and background in the sideband region (blue) and signal regions (green). In the top plot the dashed red lines denote the boundaries of the signal region defined by $m_{\gamma\gamma}$. The last bin shows overflow entries. The distributions of background events in the signal and sideband regions are seen to be similar but with visible differences, which will be learned by the interpolation aspect of the HI-SIGMA approach.
@@ 4.1 Data Smearing
The amount of smearing applied to the data is therefore a very important choice, directly impacting the sensitivity of the final analysis. In this work, we found smearing each feature by 0.2 standard deviations allowed modeling of the distributions with high accuracy while resulting in only minor sensitivity loss. The resonance feature $m_{\gamma\gamma}$, which is not modeled by the normalizing flow, is not smeared. A visualization of one feature distribution before and after smearing is shown in Fig. 2. It is possible that with an improved machine learning model, less smearing, or perhaps no smearing, would be necessary. Alternatively perhaps a less simplistic smearing procedure, perhaps smearing some features and not others, or only for data near sharp boundaries could be employed to better mitigate sensitivity loss. As such considerations are very application dependent, we leave such explorations for future work.
@@
4.2 Pre-processing, Model Architecture and Validation
@@
The generative models used to learn the signal and background conditional feature distributions, $P_{k}(\vec{x}^{\prime}|m)$, are normalizing flows. Normalizing flows Kobyzev:2021a; Rezende:2015a; Laurent:2015a estimate a given probability distribution $P_{k}(\vec{x}^{\prime}|m)$ over a feature space $\vec{x}^{\prime}$ by learning an invertible mapping between $\vec{x}^{\prime}$ and a latent $\vec{z}$ of the same dimension, $\vec{x}^{\prime}=f(\vec{z},m)$. The probability distribution can be written explicitly in terms of the function $f$, its inverse $f^{-1}$, its gradient $\nabla_{\vec{z}}f$ and the base distribution $P(\vec{z})$:
$P_{k}(\vec{x}^{\prime}|m)=P(\vec{z}=f^{-1}(\vec{x}^{\prime},m))||\det\left(\nabla_{\vec{z}}f\right)|_{\vec{z}=f^{-1}(\vec{x}^{\prime},m)}||^{-1}\,.$ (12)
$P(\vec{z})$ is chosen to be a simple base distribution that allows for sample generation and density evaluation. In this work, we consider a multivariate normal distribution $P(\vec{z})=\mathcal{N}(\vec{0},\mathbb{I})$. The function $f$ is a learnable invertible function, written in terms of Neural Networks and optimized by minimizing the negative log-likelihood derived from Eq. 12. The structure of $f$ needs to be chosen carefully to allow for evaluation not only in the forward direction ($\vec{x}^{\prime}=f(\vec{z},m)$) used for sampling but also in the backward direction ($\vec{z}=f^{-1}(\vec{x}^{\prime},m)$) necessary for likelihood evaluation and training. Additionally, the Jacobian of the function should be easily computable to ensure numerically efficient training and likelihood evaluation.
In this work, we use a Masked Autoregressive Flow papamakarios2017masked, where $f$ consists of the composition of 6 rational quadratic spline layers MPRQAF2021, each with 24 bins and a hidden size of 64. The total number of trainable parameters is 220k. BatchNorm batchnorm is used in between every flow layer. The flows are implemented using the nflows package nflows and are trained up to 100 epochs. Early stopping is triggered if no improvement is seen on the validation loss after 10 epochs. To account for training variability, five models are trained on the same dataset with different random seeds. The final model is taken as the ensemble of these 5 different models, averaging their log-probabilities to produce the final density estimate during evaluation.
The quality of the background models was checked by sampling from them and comparing the generated samples against the true background events. This is first done for background events in the sideband region in which the model was trained. A classifier trained to distinguish between the real and artificial background events, using a neural network architecture as described below, produces an AUC on of $\sim 0.501$ indicating very high quality background modeling.
We then test the extrapolation of the background models into signal region. We generate events in the signal region and compare them to true background events from the signal region. An example comparison between true events and those sampled from a single bootstrapped model is shown in Fig. 3. Despite the model never having seen events from this region, we observe quite good modeling due to the successful interpolation. There are some minor residual discrepancies in the tails which vary between the bootstraps. We train a classifier to distinguish between real and generated events in this signal region and find an AUC of $\sim 0.505$. This is slightly degraded as compared to the modeling of events in the sidebands but still quite high quality. We note that this validation is done using the entire training/validation set consisting of 700k events, which is approximately 10 times as many events as are used in the example fits in Section 5. Therefore the minor imperfections noted in this test may not be prohibitive for accurate parameter estimations on a smaller datasets with statistical uncertainties significantly larger than these minor density mismodelings. Further work would be required to properly parameterize the uncertainties on the density estimates arising from these subtle modeling imperfections.
Additionally, we note that this validation of the interpolation capabilities of the background model would not be possible in a true measurement on data, as it requires the usage of events from the signal region. However a similar validation could be performed in data in some control region, perhaps using high $M_{\gamma\gamma}$ events above the Higgs mass. A separate training could be performed on events from this region, holding out a specific mass window from the training, and then testing the model interpolation into this region.
Figure 3: Events generated from a background model (single bootstrap) as compared to true events in the signal region. To spot potential minor mismodelings, the comparison is performed using samples an order of magnitude larger than those used for the inference example. Good agreement is observed overall, with only minor discrepancies visible in the tails.
@@ 4.4 Example Shape Uncertainty
To deploy of HI-SIGMA in any real measurement, it needs to be able to incorporate systematic uncertainties. To demonstrate that this is possible, we implement a toy shape systematic within this di-Higgs example following the method detailed in Sec. 3.2. We consider a $10\%$ relative uncertainty associated with the $p_{T}$ of the $H_{bb}$ candidate, such that for a given $\nu$ we re-scale $p^{bb}_{T}\to(1+0.1\nu)p^{bb}_{T}$ mimicking the effects of a jet energy scale uncertainty. This is done for both signal and background distributions, with both shifts parameterized by a single common shared nuisance parameter. We show the effects of the $\pm 1\sigma$ variations in Fig. 4.
Figure 4: Distributions of the $p_{T}^{bb}$ feature for $\sigma$, nominal and $\sigma$ variations of the shape systematic. The other features are assumed to be unaffected.
For this simple uncertainty, where the $\nu$ dependence is certainly present but can be modelled with a low-complexity function, we consider the simple parameterization of the weights introduced in Eq. 9. We performed a closure test to verify that the interpolation works as expected. Mock datasets were generated from the $\nu=0,\pm 1$ models, and successfully fit with the full model. We verified that the fitted signal strength and nuisance parameter agreed well with the true values, $(\hat{s},\hat{\nu})=(s_{\mathrm{true}},\nu_{\mathrm{true}})$.
@@ 4.5 Comparison Analysis Strategies
The next two analysis strategies employ DNN classifiers to discriminate signal versus background, similar to what is done in many current analyses. The first strategy employs a ‘cut and fit’ approach. A DNN classifier is trained to distinguish signal versus background using the feature information, $\vec{x}^{\prime}$, but not the resonance feature $m_{\gamma\gamma}$. A cut is placed on the classifier score to reduce the background and the $m_{\gamma\gamma}$ distribution of the remaining events is fit to extract the signal. The cut value was chosen to maximize the signal significance, approximated as $s/\sqrt{b}$. Using a cut-based approach such as this is sub-optimal, as some signal events are excluded from the final analysis, and one throws away any information on signal versus background discrimination of the remaining events. Often, multiple exclusive categories based on different classifier score thresholds defined simultaneously fit, so as to mitigate this sensitivity loss. This strategy is employed in recent ATLAS and CMS analyses of di-Higgs $bb\gamma\gamma$ CMS_bbgg; ATLAS_bbgg; ATLAS_bbgg_run3. However we stick with a single category here for simplicity.
@@
The third alternate analysis method uses a classifier trained using the full event information, the features and the resonance variable, to define an optimal summary statistic. A 1D histogram of the distribution of classifier scores in data is then fit to the sum of signal and background templates to extract the signal strength. Although the classifier will learn a monotonic rescaling of the likelihood ratio, the optimal observable for signal versus background discrimination, binning will result in some loss of information. Larger numbers of bins will mitigate this information loss and result in higher sensitivity. In our example we compare two choices for the number of classifier bins, 10 and 50.
Both DNN-based analyses use almost the same DNN architecture, differing only on the size of the input. The input layer of both DNNs has a BatchNorm layer batchnorm to scale the features appropriately. Both DNNs possess five inner layers with 32,64,64,32 and 8 neurons, respectively, and a use a and a SiLU activation function after the first layer and Leaky ReLU for the other layers. The final layer consists of a single single neuron and a sigmoid activation function. To avoid overfitting, a Dropout layer is added after the input and inner linear layers, with a dropout probability of $0.2$. Both models are implemented in Pytorch and trained to minimize the Binary Cross-Entropy loss, with the same early stopping procedure as the normalizing flows but with $15$ epochs. The DNN architecture was not extensively optimized, but it was checked that increasing the size of the hidden layers did not improve performance.
@@
This third alternate analysis requires an accurate model of the background classifier score shape. Because the classifier score distribution has complex multivariate correlations with all the input features, estimating the background shape using data-driven methods can be difficult and methods are usually analysis-specific. For purposes of sensitivity comparison, in this study we consider an idealized setting where the signal and background classifier score shapes are known exactly. This allows a comparison with the HI-SIGMA sensitivity in an idealized setting in which systematic uncertainties are not assessed on either method. However, it should be noted that the HI-SIGMA method automatically includes a data-driven background estimate, whereas for the classifier score analysis such an estimate may be quite difficult.
@@
4.6 Toy Simulation Mismodeling Example
We should note that the optimality of the DNN-based analyses depends on the accuracy of the background simulations used for training, which is the very issue that motivates HI-SIGMA. Even if a subsequent data-driven background estimate is performed, the performance of the DNN classifier may be sub-optimal if the background simulations used for its training are inaccurate. Since HI-SIGMA learns the background likelihood directly from data, it does not suffer from this limitation and may offer an advantage over the DNN in this case.
To study this potential effect, we assess the performance of HI-SIGMA and the comparison strategies under data-MC mismodeling. We mimic data-MC differences by slightly distorting the the background ’simulation’ sample used to train the DNN classifier. We then use these classifiers to perform inference on the original data, with no distortion. For the binned DNN score fit, we still assume the shape of the background DNN score shape is known exactly from some data-driven method. Therefore, inference with this distortion-trained classifier is not biased, just sub-optimal. In contrast, HI-SIGMA learns the background density directly from the data, and is therefore unaffected by such background ‘simulation’ mismodelings.
The distortions applied in our mock study, shown in Fig. 5, are as follows. The widths of the $\Delta R_{bb}$ and $\Delta R_{\gamma\gamma}$ distributions are increased by $\sim$10%. The Hbb $p_{T}$ and Hγγ $p_{T}/m$ distributions have their means shifted by 5% and are given 3% additional Gaussian random noise.
Figure 5: Illustrations of the original (blue) and distorted (green) feature distributions of the background sample. The distorted sample is used to train classifiers, approximating data-MC modeling imperfections. See text for details.
5 Results
The different analysis methods detailed in Section 4 are evaluated on 5 independent data samples, each containing 40,000 events. Unless otherwise noted, for each data sample 200 signal events are injected, leading to an approximate significance with the mass-only fit of ${\sim}3\sigma$. The different analysis methods are then applied to these datasets and compared. Examples of the mass only fit, classifier cut and fit, and binned DNN score fit are shown in Fig. 6.
Figure 6: Examples of the mass only fit (upper left), cut+fit approach (upper left) and classifier score fit with 10 (lower left) and 50 (lower right) bins.
5.1 Visualization of the HI-SIGMA Fit
The HI-SIGMA fit across the multi-dimensional space can also be visualized. To do this, we sample events from the signal and background models to construct ‘templates’ of our signal and background distributions. These templates are then scaled by the best fit normalizations and compared to the observed data events. This comparison can be done in any way that is desired. The simplest choice is to consider 1D projections of each feature, which we visualize as histograms. Examples are shown in Fig. 7. We stress that the binning is for visualization purposes only, as the actual model and fit are unbinned. For each feature one can see an excess of data events above the background model which is well-fit by the signal template.
@@ 5.1 Visualization of the HI-SIGMA Fit
While the visualization in Fig. 7 is useful, these 1D projections do not ensure the data is being well described across all regions of the multi-dimensional space. Fortunately, because HI-SIGMA includes a full multi-dimensional model of the signal and background densities, one can check the fit agreement in sub-regions of the space easily. We consider a check of the most signal-sensitive region of the space, defined by placing a cut on the DNN classifier score >0.9>0.9 and restricting $m_{\gamma\gamma}$ to the signal region, 115–135 GeV. Placing these cuts on the events generated from our model as well as the data allows us to check modeling in this sub-region of the space. Such a comparison is shown in Fig. 8. We see that in this sub-region the data is well described by the model and signal contribution is much more visible. We stress that such a visualization can be performed for any sub-region of the feature space $\vec{x}^{\prime}$, but cannot be done for any region defined by an observable outside the feature space of the model (e.g. checking modeling in different regions of $\eta$ for our example). Although this method is helpful in assessing the quality of the fit in different regions of the multi-dimensional space, a full assessment of the goodness-of-fit of an unbinned high-dimensional fit would require further study Williams:2010vh.
Figure 8: Visualization of a subspace of the HI-SIGMA fit, the most signal-sensitive region, defined by a requirement on the supervised classifier score >0.9>0.9.
5.2 Performance Quantification
The average uncertainty on the signal strength parameter (defined as $\frac{s}{s_{\mathrm{true}}}$) from the fit to the 5 independent datasets is summarized in Fig. 9. The HI-SIGMA approach is shown to improve upon the sensitivity of the cut+fit approach by $\sim 40\%$, and by $\sim 20\%$ as compared to the classifier score fit with 10 bins. The performance of HI-SIGMA is similar to the classifier score fit with 50 bins. We stress that this is an ‘idealized’ comparison, in which no systematic uncertainties have been included for any of the analysis strategies, and strategies for realistic estimates of the background in the classifier score fit have not been considered. The different analysis strategies would necessitate different systematic uncertainties, which may change their relative sensitivities. However, the ability of HI-SIGMA to match an idealized DNN score fit with a large number of bins (50), while including a realistic background estimation strategy, demonstrates the potential of the approach.